OpenStudy (anonymous):

4^x - 2^x = 3 Solution? Is it unsolvable?

OpenStudy (anonymous):

It is solvable ,

OpenStudy (anonymous):

Can you show the source of the question ( snapshot ) ?

Parth (parthkohli):

\[\implies \left(2^x\right)^2 - (2^x) = 3\]

Parth (parthkohli):

\[\implies (2^x)^2 - (2^x) = 3\]\[\implies (2^x)^2 - (2^x) - 3 =0\]Find the solutions of this quadratic. Ignore the negative solutions as \(2^x\) cannot be negative.

OpenStudy (anonymous):

Tried using the quadratic formula. I ended up with (2^x +- 2^x * sqrt(3)) / 4^x = 0 and I cannot seem to get anywhere else from that. Also I have tried entering this into wolfram/mathway and it gives no solution or impossible solutions.

OpenStudy (asnaseer):

If you use the hints provided by @ParthKohli then you should find a valid solution. It might help if you substituted: \(y=2^x\) and solved for \(y\) first.